OpenStudy (anonymous):

1/2_3/7 ? A. < B. > C.=

hero (hero):

\[\frac{1}{2} \text{___} \frac{3}{7}\] Cross Multiply: 1(7) ___ (2)(3) 7 > 6 Therefore: \[\frac{1}{2} > \frac{3}{7}\]

OpenStudy (anonymous):

Thanks!

hero (hero):

yw

hero (hero):

I think you'll be able to do these now. All you have to do is just cross multiply to figure out greater than, less than, or equal.